The network below shows a network of railway lines from city 1 to city 10. The numbers on the arcs represent the number of cars per hour that can pass along that segment of track 400 600 300 150 500 200 400 600 450 10 6 300 300 300 200 300 500 500. 400 350 Find the maximal flow in cars per hour that can pass through the network from node 1 to node 10. In each blank, enter the following information. If there is no flow along an arc, enter 0. maximal flow (enter as 1400, blank #1 not 1,400) flow from node 1 to node 2 blank #2 blank #3 flow from node 1 to node 3 flow from node 1 to node 4 blank #4 flow from node 2 to node 5 blank #5 flow from node 2 to node 6 blank #6 flow from node 3 to node 7 flow from node 3 to node 9 blank #7 blank #8 blank #9 flow from node 4 to node 7 blank #10 flow from node 5 to node 8
flow from node 5 to node 8 blank #10 flow from node 5 to node 10 blank #11 flow from node 6 to node 5 blank #12 blank #13 flow from node 7 to node 3 blank #14 flow from node 7 to node 9 blank #15 flow from node 8 to node 5 blank #16 flow from node 8 to node 9 blank #17 flow from node 8 to node 10 blank #18 flow from node 9 to node 8 blank #19 flow from node 9 to node 10
Answer #1

Answer:---------------

Min XF1
s.t.

X12 + X15 + X16 - XF1= 0

X23 + X24 - X12= 0

X34 - X23= 0

X48 + X4F - X24- - X34 - X84= 0

X57 - X15= 0

X67+ X69 - X16 - X76= 0

X76+ X79 - X57 - X67= 0

X84+ X89 + X8F - X48 - X98= 0

X98+ X9F - X69 - X79 - X89= 0

XF1 -X4F - X8F -X9F= 0

X12  ≥  500

X15300

X16600

X23300

X24400

X34150

X48400

X4F600

X57400

X67300

X69500

X76200

X79350

X84200

X89300

X8F450

X98300

X9F500

Xij0 for all i,j